OpenStudy (anonymous):

I will give medal and a fan to the person who helps me ....pretty please !!!!! Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power the cube root of 5 squared the twelfth root of 5 the square root of 5 the cube root of 5 the fourth power

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (kewlgeek555):

Rational Exponent Property: Fractional Powers, or where a number is raised to a fraction, can be converted to a radical. The numerator becomes the exponent, and the denominator becomes the index of the radical. \[x ^{\frac{ 1 }{ n }}=\sqrt[n]{x}\]\[x ^{\frac{ m }{ n }}=\sqrt[n]{x^m}\]\[(x \neq 0)\]

OpenStudy (anonymous):

oh that makes so much more sense, i can figure it out now thanks so much for you help @kewlgeek555

OpenStudy (anonymous):

you get a medal

OpenStudy (kewlgeek555):

Yes. Remember to memorize all the properties of exponents for Module Four. It will be helpful. Especially since you will be working with a lot of functions and equations. :)

OpenStudy (kewlgeek555):

And thank you. ;D

OpenStudy (anonymous):

i will :) thanks for the medal